GreenHab Report – November 13th

Crew 215 GreenHab Report 13-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation 12 midday then reclosed 5pm.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: 20 litres (2.5 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank – tank was completely filled during afternoon.
Time(s) of watering for crops: 9am, 3pm.
Changes to crops: 2 tomato species, 2 spinach species, rocket planted, rotted Aloe Vera plants removed. The Aloe Vera were most likely over watered, these species preferring a dry sandy soil. The had also some frost damage when delivered.

Established plants introduced to Greenhab 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
[2x] small Aloe Vera, in medium metal planter, frost damaged (removed 1311219)

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)

Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted in Greenhab 10/11/2019 (Large metal planting tub)

Corinader – Coriandrum Sativum

Seeds Planted in Greenhab 12/11/2019 (2 smaller planting tubs)

Carrots ‘Purple Sun F1’
Radish ‘icicle’

Seeds Planted in Greenhab 13/11/2019
, in medium metal planter

Other seeds stored in the Greenhab are identified as –

Journalist Report – November 13th

Wed 13 Nov Sol 3
by Guy Murphy
Today was Jennifer Lane’s birthday on Mars. While she was out on EVA during the afternoon with Andrew and Shane, the upper level of the Hab was decorated with balloons and streamers. Larissa baked a sponge cake with cherry cream, with Jennifers name spelt on the top in coloured icing balls. For main course, a large focaccia bread topped with olives was also created to have with soup. Our festive dinner was celebrated with a toast of bubbly (carbonated water).
Celebrating special occasions such as this on long duration missions is a tradition in places such as Antarctica. They offer an opportunity to vary an otherwise repetitive weekly schedule, when harsh weather conditions confine crew indoors. Crew may pack special foods, drink and amusements ahead of time to surprise the other team members on the day, as we did this evening. They can also help bring together international crew, where special events can bring together the whole crew, providing an opportunity to share national cuisines. The only non-Australian in Crew 215 is Larissa Wilson, who is from New Zealand. She had already tried Vegemite.
Tomato, spinach and rocket seedlings were planted in the Greenhab today, hopefully to gift fresh green garden salads to future crew.

EVA Report – November 13th

Crew 215 EVA Report 13 Nov 2019
EVA # 03
Author: Jennifer Lane
Purpose of EVA: Scouting for and collecting gypsum for use in creating Martian concrete
Start time: 0200
End time: 1555
Narrative: Heading north on Cow Dung road, we stopped at White Moon and immediately found our desired gypsum. We collected a bucket full in individual zip lock bags. These will be crushed over the weekend (while we’re unable to conduct EVAs due to a solar event) and mixed with an aggregate to create concrete; testing the ability for in-situ recourse utilization on Mars. On our return journey, we began having issues with the Opportunity Rover. Lucky we were close enough to have a visual on the hab, and so walked the remaining distance before contacting outpost for assistance with retrieving the vehicle.
Destination: White Moon
Coordinates (use UTM NAD27 CONUS): 517500E 4254650N
Participants: Andrew Wheeler, Shane Usher and Jennifer Lane
Road(s) and routes per MDRS Map: Cow Dung Road onto Brahe Hwy 1572
Mode of travel: Curiosity and Opportunity Rover

Research Report – November 13th

Science Report 13 November 2019
SOL: 3
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: N/A
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The third EVA for science saw us returning to the White Moon location at the junction of Brahe Hwy and Cow dung Road (517790E 4254750N) to collect more gypsum samples for ISRU cement experiments. It was noted that, at this location out of sim, at least one vehicle has been driving through the exact locations of our previous sampling efforts. Fortunately, this has no effect on our sampling, but it does suggest that our activities are been monitored extremely closely. See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin monitoring vests. In addition, equipment for the retrieval of an injured marstronaut was inventoried prior to a rehearsal for an actual event. Subject to approval, a simulated emergency retrieval of the engineer will be held 14Nov2019 at the conclusion of his engineering EVA.
Additional activities: Nil.

Operations Report – November 13th

Crew 215 Operations Report 13-Nov-2019
SOL: 3
Name of person filing report: Shane Usher
Non-nominal systems: Opportunity, Curiosity
Notes on non-nominal systems: Suit 2 battery replaced. Opportunity broke down.
Curiosity plugged in with light flashing, but not charging battery.
Generator: run
Hours run: 14h 00 min
From what time last night: approx. 1745
To what time this morning: 0745
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 72%
Diesel Reading – 50% –
Station Propane Reading – 81%.
Ethanol Free Gasoline: N/A
Water (loft tank): 47 gallons
Water Meter: 1457711 units
Water (static tank): 420
Static to Loft Pump used – yes
Water in Green Hab: 287.8 gallons, 20 L (5.7 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 110.5
Beginning charge: (Before EVA): 62%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.
Opportunity rover used: yes
Hours: 66.2
Beginning charge: 100%
Ending charge: 55%
Currently charging: yes
Curiosity rover used: yes
Hours: 116.7
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 67%
Currently charging: yes
Notes on rovers: Opportunity broke down. Curiosity does not appear to be charging,
light is flashing but % charge is not increasing.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional, but there was a period of no internet access.
Summary of suits and radios: Suit 2 battery replaced, Suit 5 Helmet photo sent.
Summary of Hab operations: No new issues.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried
out during the day, plants were watered (20L). The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity now may not be usable.
Waiting to see if Curiosity is charging.

Sol Summary – November 13th

Crew 215 Sol Summary Report 13-NOV-2019
Sol: 3
Summary Title: A birthday on Mars.
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: Sol 3 had Jennifer wake to a Martian birthday. After a special pancake breakfast, activity in the GreenHab occupied our attention. During this time we were advised of a possible G1 storm the might affect us on the 16/17th. Though we were being filmed and videoed for publicity, we continued with our data collection and preparations for experiments. Equipment designed to retrieve a disabled marstronaut was laid out in preparation for practice. If approved, we will be working through this scenario at the end of the engineering EVA tomorrow. The afternoon saw us returning to the White Moon location (517724E 4254500N) at the junction of Brahe Hwy and Cow Dung
Road where we retrieved gypsum for further experimentation. Unfortunately, our prior activities had been closely monitored and, on arrival, we found that wheel tracks had cut through our previous sample locations and across our boot prints. On return to the hab, Opportunity rover displayed a warning light and cut power 200m from the hab, meaning that the EVA ended on foot. This evening saw a special dinner of focaccia, soup and birthday cake for Jennifer.
Look Ahead Plan: Planning an EVA to MM5 (518462E 4248949N) to resume collection of micrometeorite samples and a demonstration of returning a disable marstronaut into the hab during the afternoon engineering EVA. See EVA requests. Continue monitoring of psycho-social stressors.
Anomalies in work: Opportunity depowered during EVA return to hab.
Weather: Marginally below freezing overnight, light winds, clear skies.
Crew Physical Status: Healthy
EVA: Gypsum sampling EVA. See EVA report
Reports to be file: Sol Summary Report, Operations Report, Science Report, GreenHab
Report, EVA Request, EVA Report, Journalist Report, Daily Photos.
Support Requested: N/A

Journalist Report – November 14th

Thurs 14 Nov Sol 4
by Guy Murphy
This morning I undertook my first extravehicular activity (EVA) of this Crew rotation.
Steve, Larissa and I took a pair of small 4 wheeled electric rovers out along the road
to the south on the campus where we are sampling for micro-meteorites. The simulated
EVA space suits restrict the line of sight, requiring the wearer to orientate their
body to look to the side rather than turn their head. This is most noticeable driving
the rovers, as you cannot easily see behind you from the driver’s seat. The suits
include gloves, which remove the sense of touch that would normally inform contact with
surfaces and objects around you. The enclosed helmets create a void of silence, verbal
contact only possible using radios.
We spent an hour and a half at the sample collection site. Radio dialogue with the
others was supplemented by hand signals. Space suits can get hot when you are working,
despite internal air conditioning and cooling. If your face gets itchy, you cannot
scratch your nose. But its very beautiful being outside in this landscape and an
opportunity for some physical exercise. Working together we completed nearby 1/3rd of
the mission’s planned sampling at this site before returning to the Hab.
On Mars, humans are inherently more efficient than robots when it comes to field
research, and engineering requiring human cognition and dexterity. This is why we need
humans there. EVAs will only be done when strictly necessary, as any venture outside is
inherently dangerous. Exposure to background radiation will discourage individuals from
logging up too much EVA time. People will never go out by themselves, only in pairs or
more. Living indoors for many months on end, any chance to be outside under open skies
will be highly valued.
A restricted Martian pantry need not result in a bland diet. This morning Larissa
served freshly baked croissants, produced with dried butter and flour. We have been
making blocks of hard cheese by filling tins with dried grated cheese, pouring in hot
water, pressing the cheese with forks, then draining and refrigerating. Once taken from
their moulds and aired, the sliced cheese is great to serve on dry biscuits with pickles.

Copyright © The Mars Society. All rights reserved. | Main Site